Phantomdex:

There is a 58% chance that a painter will work today, a 27% chance that it will storm today, and a 13% chance of both happening. Determine the probability that it will storm today or the painter will work today. 16% 31% 72% 85%

toga:

To determine the probability that it will storm today or the painter will work today, we can use the formula: P(A or B) = P(A) + P(B) - P(A and B) Where A represents the event that the painter will work today, and B represents the event that it will storm today. Given that P(A) = 0.58, P(B) = 0.27, and P(A and B) = 0.13, we can substitute these values into the formula: P(A or B) = 0.58 + 0.27 - 0.13 = 0.72 Therefore, the probability that it will storm today or the painter will work today is 72%, which is option C.